the sum of two numbers is 12 the difference of the same two numbers is negative four what are the numbers?
Answer by oberobic(2304)

x + y = 12
so
x = 12 -y
.
x - y = -4
so
x = y-4
.
x=x
so
12-y = y -4
add y to both sides
12 = 2y -4
add 4 to both sides
16 = 2y
2y = 16
divide both sides by 2
y = 8
.
x + y = 12
x + 8 = 12
x = 4
.
Check using the other equation.
.
x - y = ?
4 -8 = -4
Correct.
.
Answer: y=8 and x=4
.
Done.